# I2S-TFCKD: Intra-Inter Set Knowledge Distillation with Time-Frequency Calibration for Speech Enhancement

### v1

@@ V-A3 Analysis of Evaluation Metrics
Table III presents the comparison of objective speech metrics across different distillation strategies on the multi-SNR test set, and statistical significance analysis is conducted in the average metric comparison under three SNRs. We use the significance level $p$ to measure the statistical improvement of each distillation method compared to the student model. From the results, all distillation methods improve the student’s objective metrics across the three SNR levels, narrowing the gap between the teacher and student. Among them, M1 that performs layer-wise distillation, underperforms other methods, showing no significant improvement ($p>0.1$) over the student model in any of the six metrics. In contrast, M2 that extends single-layer distillation to attention-based multi-layer distillation, outperforms M1 across all metrics and achieves secondarily significant improvements ($0.10.05<p<$ 0.1) in CSIG and STOI compared to the student model. This indicates that cross-layer interactions provide additional representational knowledge. Building on this, M3 that introduces time-frequency cross-calibration, achieves further metric advantages, with statistically significant improvements ($p<0.05$) in CSIG, COVL, and STOI compared to the student model. This demonstrates that cross-calibration captures discriminative information across frequency bands and temporal segments at different layers, enabling more refined teacher-student layer matching. M4 that introduces residual fusion for inter-set distillation, achieves the best metric results among all strategies, with statistically significant improvements ($p<0.05$) in five metrics under average SNR compared to the student, except for SISNR. This indicates that collaborative distillation within and between correlated sets effectively promotes global knowledge circulation, significantly improving the enhancement effect of the student model.
V-B Comparative Experiments on the DNS Benchmark
TABLE IV: Comparison of objective speech evaluation metrics between distilled and undistilled models on the DNS test set.
Methods Casual & NFI Param. (M) FLOPs (G/s) PESQ STOI SISNR (dB)
Noisy - - - 1.58 0.915 9.07
RNNoise [34$\dagger$1 0.1 0.04 1.97 - -
NSNet [35$\dagger$ 1.3 0.13 2.15 0.945 15.61
DTLN [36$\dagger$ 1.0 0.25 - 0.948 16.34
DCCRN [37$\dagger$ 3.7 14.37 2.80 0.964 16.38
FullSubNet+ [38$\dagger$ 8.7 31.12 2.98 0.967 18.34
FRCRN [6$\dagger$ 6.9 12.30 3.23 0.977 19.78
CTS-Net [39$\dagger$ 4.4 5.57 2.94 0.967 17.99
PrimeK-Net [40$\ddagger$ 1.4 44.64 3.34 0.978 19.23
@@ V-B Comparative Experiments on the DNS Benchmark
-MSE [32$\ddagger$ 0.6 2.44 2.82 0.962 16.60
-CLSKD [26$\dagger$ 0.6 2.44 2.84 0.963 17.08
-UCLFWPKD [16$\dagger$ 0.6 2.44 2.94 0.966 17.16
-ABC-KD [14$\ddagger$ 0.6 2.44 2.90 0.964 17.14
-Two-Step KD [15$\ddagger$ 0.6 2.44 2.93 0.965 17.05
-M1 (FitNet [33$\ddagger$) 0.6 2.44 2.83 0.962 17.22
-M2 (SemCKD [30$\ddagger$) 0.6 2.44 2.89 0.965 17.36
-M3 (TFCKD) 0.6 2.44 2.97 0.967 17.58
-M4 (I2S-TFCKD) 0.6 2.44 3.03 0.968 17.74
@@
Fair comparisons of various algorithms are conducted on the DNS non-reverb test set. Table IV presents the objective speech quality evaluation metrics of various SE methods. We provide details on the causality of each model and whether it uses future frame information (Casual and No Future Information, Casual & NFI). Unlike the 40 ms look-ahead constraint of the DNS challenge, we impose strict restrictions on both the causal architecture and the exclusion of future information to better adapt to real-time applications. Additionally, the model parameter (Param.) count and floating-point operations (FLOPs) per second are provided to evaluate deployment feasibility on edge devices. Among the SE models in Table IV, NSNet, RNNoise, and DTLN are low-complexity real-time SE algorithms primarily based on LSTM networks. These models are fully causal and do not utilize future information. In contrast, DCCRN, FullSubNet+, FRCRN, CTS-Net, and PrimeK-Net fail to satisfy both causality and NFI constraints. Although some architectures, such as DCCRN and CTS-Net, employ causal designs, they still leverage partial future frame information during training and inference. Our distillation baseline DPDCRN, composed of causal convolutions and GRU layers, is strictly causal and operates without any future frame information.
V-B1 Analysis of Distillation Frameworks
Figure 6: The scatter-bubble distribution of PESQ, FLOPs, and model parameters for SE models on the DNS test set.
We compare the effects of different distillation frameworks applied to the base model DPDCRN in Table IV. Among them, MSE that directly narrows the output distance, only achieves slightly higher results than the student model on PESQ and STOI metrics and even falls below the student on the SISNR metric. In contrast, M1 achieves effective improvements compared to output-level distillation through layer-wise feature distillation. On this basis, both CLSKD and M2 incorporate multi-layer knowledge integration to enable cross-layer interaction between the teacher and the student, thus achieving more considerable gains in PESQ and STOI. The distillation advantages of the above multi-layer knowledge transfer frameworks demonstrate that the intermediate representations of the model contain rich transferable information. Recent studies such as UCLFWPKD, ABC-KD, and Two-Step KD design a targeted cross-layer distillation framework according to the characteristics of the SE task, achieving further improvements. In this paper, M3 introduces time-frequency cross-calibration based on the multi-layer distillation framework of M2, achieving notable improvements across all three metrics. This highlights that computing multi-layer calibration weights along both the temporal and spectral domains enables more effective knowledge allocation. Extending M3 to the intra-inter set distillation framework (M4, I2S-TFCKD) achieves the best performance among all distillation methods, demonstrating that knowledge circulation within and between correlated sets further facilitate teacher-student knowledge transfer.
V-B2 Comparison with State-of-the-art SE Methods
@@
Table IV provides the metric performances of current SOTA SE algorithms on the DNS test set. Additionally, Fig. 6 shows the scatter-bubble distribution of PESQ with model parameters and FLOPs. It can be observed that current real-time SE algorithms such as RNNoise, NSNet, and DTLN, despite having extremely low parameters and FLOPs, still exhibit a large performance gap compared to high-complexity SE models. Models like FRCRN and FullSubNet+ have metric advantages, but their high parameter counts (>5M) and FLOPs (>10G) restrict edge-side applicability. The recently proposed SE model PrimeK-Net achieves substantial parameter reduction through targeted architectural design, but its 44.64G FLOPs still impose enormous deployment pressure. This paper applies intra-inter set distillation with time-frequency calibration (I2S-TFCKD) to the student model DPDCRN-S, attaining competitive results across three metrics with low parameter count (0.6M) and computational complexity (2.4G FLOPs).
V-C Comparative Experiments on the L3DAS23 Benchmark
V-C1 Boxplot Analysis of Distillation Strategies
@@
Fig. 7 presents a boxplot comparison of representative metric results (PESQ and T1 Metric) for different distillation frameworks on the L3DAS23 validation and development sets. From the perspective of the mean value of the metrics, MSE still remains inferior to feature-based distillation strategies, indicating that the utilization of intermediate representation information is also crucial for multi-channel SE distillation. Among all distillation algorithms, the proposed I2S-TFCKD method achieves the highest mean values across both metrics. Compared with the second-highest performing UCLFWPKD, the proposed framework has advantages in both the average metric and the concentration degree of high values, further validating the positive contributions of intra-inter set knowledge flow and time-frequency cross-calibration to the overall distillation effect.
V-C2 Metric comparison Across Models
Table V presents objective metric results on the L3DAS23 development set, comparing multi-channel SE models with and without KD. The base model DPDCRN achieved first place on the blind test set of the L3DAS23 challenge and is causally adjusted in this paper. The teacher backbone DPDCRN-T demonstrates substantial metric advantages over the causal framework EaBNet across all metrics, and ranks second only to the recently proposed non-causal Spatial Net, but the latter still maintains extremely high FLOPs. The compressed student DPDCRN-S exhibits performance degradation compared to the teacher, particularly in PESQ and the WER indicator. Application of various distillation strategies to the student results in varying degrees of improvement across metrics. The proposed I2S-TFCKD method achieves optimal distillation effect, with indicator gains of 0.218 in PESQ, 0.035 in WER, 0.02 in STOI, and 0.028 in the T1 Metric compared to the undistilled student. Overall, the distilled student model maintains causal inference with low computational overhead while achieving competitive performance compared to SOTA SE methods in enhancement results.
@@ V-C2 Metric comparison Across Models
TABLE V: Comparison of objective speech evaluation metrics between distilled and undistilled models on the development set in the SE track of L3DAS23.
@@
Neural Beamforming [41$\dagger$ 5.5 32.15 - 0.569 0.684 0.557
EaBNet [42$\dagger$ 2.8 7.38 - 0.549 0.724 0.587
CCA Speech [43$\dagger$ - - - 0.293 0.836 0.771
DeFT-AN [44$\ddagger$ 2.7 37.99 1.976 0.137 0.866 0.864
Spatial Net [45$\ddagger$ 1.6 46.30 2.264 0.133 0.910 0.889
@@
-MSE [32$\ddagger$ 0.6 2.44 1.759 0.181 0.847 0.833
-CLSKD [26$\dagger$ 0.6 2.44 1.768 0.178 0.851 0.837
-UCLFWPKD [16$\dagger$ 0.6 2.44 1.852 0.162 0.861 0.850
-ABC-KD [14$\ddagger$ 0.6 2.44 1.795 0.169 0.856 0.843
-Two-Step KD [15$\ddagger$ 0.6 2.44 1.832 0.169 0.859 0.845
-M1 (FitNet [33$\ddagger$) 0.6 2.44 1.774 0.179 0.850 0.836
-M2 (SemCKD [30$\ddagger$) 0.6 2.44 1.851 0.161 0.859 0.849
-M3 (TFCKD) 0.6 2.44 1.914 0.152 0.866 0.857
-M4 (I2S-TFCKD) 0.6 2.44 1.929 0.150 0.870 0.860
@@
VI Conclusion
In this paper, we proposed an intra-inter set distillation framework with time-frequency calibration (I2S-TFCKD) for SE model compression. Specifically, I2S-TFCKD partitions intermediate layers of SE models into multiple correlated sets. Within each set, intra-set cross-layer knowledge transfer is performed, and representative features are generated through residual fusion. The representative features of multiple correlated sets are then integrated into a fused feature set to enable inter-set distillation. Furthermore, we designed a multi-layer time-frequency cross-calibration distillation tailored to SE tasks: the temporal stream captures importance distributions across speech frames, while the spectral stream focuses on distillation contributions of layer-wise spectral features. We verified the effectiveness of time-frequency calibration and intra-inter set distillation through ablation studies. Experimental results on two public datasets (DNS and L3DAS23) demonstrated that the proposed distillation framework stably and effectively improves the enhancement performance of the low-complexity student model, narrows the gap with the teacher and outperforms other distillation strategies across various metrics. Future research will focus on multi-teacher distillation for SE models, enabling an elite student to handle various acoustic scenarios through knowledge integration from multiple teachers.

### v2

@@ 5.2.2 Comparison with state-of-the-art SE methods
[…]
@@
[…]
-M1 (Layer-wise KD) 0.6 2.44 2.83 0.962 17.22
[…]
-M3 (Intra-set matching with TFCKD) 0.6 2.44 2.97 0.967 17.58
-M4 (I2RF-TFCKD) 0.6 2.44 3.03 0.968 17.74
@@
Table 4 provides the metric performances of current SOTA SE algorithms on the DNS test set. Additionally, Fig. 7 shows the scatter-bubble distribution of three metrics (PESQ, STOI, and SISNR) with model parameters and FLOPs. It can be observed that current real-time SE algorithms such as RNNoise, NSNet, and DTLN, despite having extremely low parameters and FLOPs, still exhibit a large performance gap compared to high-complexity SE models. Models like FRCRN and FullSubNet+ have metric advantages, but their high parameter counts (>5M) and FLOPs (>10G) restrict edge-side applicability. The recently proposed SE model PrimeK-Net achieves substantial parameter reduction through targeted architectural design, but its 44.64G FLOPs still impose enormous deployment pressure. This paper applies intra-inter set distillation with time-frequency calibration (I2RF-TFCKD) to the student model DPDCRN-S, attaining competitive results across three metrics with low parameter count (0.6M) and computational complexity (2.4G FLOPs).
5.3 Comparative experiments on the L3DAS23 benchmark
5.3.1 Boxplot analysis of distillation strategies
@@
Fig. 8 presents a boxplot comparison of representative metric results (PESQ and T1 Metric) for different distillation frameworks on the L3DAS23 validation and development sets. From the perspective of the mean value of the metrics, MSE still remains inferior to feature-based distillation strategies, indicating that the utilization of intermediate representation information is also crucial for multi-channel SE distillation. Among all distillation algorithms, the proposed I2RF-TFCKD method achieves the highest mean values across both metrics. Compared with the second-highest performing UCLFWPKD, the proposed framework has advantages in both the average metric and the concentration degree of high values, further validating the positive contributions of intra-inter set knowledge flow and time-frequency cross-calibration to the overall distillation effect.
5.3.2 Metric comparison across models
@@
Table 5: Comparison of objective speech evaluation metrics between distilled and undistilled models on the development set in the SE track of L3DAS23.
@@
[…]
@@
[…]
-M1 (Layer-wise KD) 0.6 2.44 1.774 0.179 0.850 0.836
[…]
-M3 (Intra-set matching with TFCKD) 0.6 2.44 1.914 0.152 0.866 0.857
-M4 (I2RF-TFCKD) 0.6 2.44 1.929 0.150 0.870 0.860
@@
Table 5 presents objective metric results on the L3DAS23 development set, comparing multi-channel SE models with and without KD. The base model DPDCRN achieved first place on the blind test set of the L3DAS23 challenge and is causally adjusted in this paper. The teacher backbone DPDCRN-T demonstrates substantial metric advantages over the causal framework EaBNet across all metrics, and ranks second only to the recently proposed non-causal Spatial Net, but the latter still maintains extremely high FLOPs. The compressed student DPDCRN-S exhibits performance degradation compared to the teacher, particularly in PESQ and the WER indicator. Application of various distillation strategies to the student results in varying degrees of improvement across metrics. The proposed I2RF-TFCKD method achieves optimal distillation effect, with indicator gains of 0.218 in PESQ, 0.035 in WER, 0.02 in STOI, and 0.028 in the T1 Metric compared to the undistilled student. Overall, the distilled student model maintains causal inference with low computational overhead while achieving competitive performance compared to SOTA SE methods in enhancement results.
6 Conclusion
In this paper, we proposed an intra-inter set distillation framework with time-frequency calibration (I2RF-TFCKD) for SE model compression. Specifically, I2RF-TFCKD partitions intermediate layers of SE models into multiple correlated sets. Within each set, intra-set cross-layer knowledge transfer is performed, and representative features are generated through residual fusion. The representative features of multiple correlated sets are then integrated into a fused feature set to enable inter-set distillation. Furthermore, we designed a multi-layer time-frequency cross-calibration distillation tailored to SE tasks: the temporal stream captures importance distributions across speech frames, while the spectral stream focuses on distillation contributions of layer-wise spectral features. We verified the effectiveness of time-frequency calibration and intra-inter set distillation through ablation studies. Experimental results on two public datasets (DNS and L3DAS23) demonstrated that the proposed distillation framework stably and effectively improves the enhancement performance of the low-complexity student model, narrows the gap with the teacher and outperforms other distillation strategies across various metrics.
Despite the progress achieved by our distillation method in the compression of SE models, there are still some limitations that need to be addressed. For example, the performance of our method on tasks such as speech separation and echo cancellation requires further validation. In addition, it remains unclear whether our method can be extended to multi-teacher distillation scenarios. In future work, we will continue to explore the application of the proposed method to more diverse speech signal processing tasks, and further investigate multi-teacher distillation for speech models.
7 Acknowledgments
This work was supported in part by the Natural Science Foundation of the Jiangsu Higher Education Institutions of China under Grant No. 25KJB510010, and the National Natural Science Foundation of China under Grant No. 62472227 and No. 61871213.